# Cell bulging and extrusion in a three-dimensional bubbly vertex model for curved epithelial sheets

### v1

@@ V Discussion
Here we have introduced the bubbly vertex model (BVM) for multicellular epithelial monolayers as a generalization of the standard vertex model (VM). In contrast to the VM, the BVM allows for curved interfaces, which are known to occur in flat epithelia and even more so in curved ones. Because curved epithelial sheets naturally have topological defects due to Euler’s polyhedron theorem, we asked if cell bulging and extrusion might occur naturally at such defects. By comparing the BVM with the VM, we found that allowing for cell interface curvature dramatically lowers the conical buckling threshold around such defects.
@@
In order to better understand the physical mechanism behind this observation, we developed a theory of a discrete defect cell in a mean-field tissue. This approach provides a mechanistic model for cell extrusion and it suggests that lowering the apico-basal tension in the cell to be extruded, and thus lowering the contractile forces and the mechanical stability of its cortex, provides a mechanism to extrude cells. Line tensions, as they could arise from supracellular actin structures, facilitate extrusion, but do not allow for the extrusion of partially extruded or fully bulged cells as they are observed in some experimental cases. Luminal pressures and interfacial rigidities directly reduce the influence of the topological defect on cell shape, which could serve as a mechanism for symmetry breaking in systems such as organoids.
Figure 10: Cell bulging as it is observed in mouse small intestine organoids and in the bubbly vertex model. (a) Organoid in which a single cell bulges outward into a partially extruded state. (b) Slice through a simulation of a spherical organoid with the bubbly vertex model with an apico-basal tension of $\Gamma=0.9$ and a pentagonal defect tension of $\Gamma_{\mathrm{defect}}=0.65$, where a cell is partially extruded. (c) Organoid developing a bubbly shape with varying cellular opening angles (arrows). (d) Slice through a simulation of a spherical shell in the bubbly vertex model with $\Gamma=0.9$. In (a,c) green is membrane ($\beta$-catenin), red is F-actin and blue nuclei. Microscopy data courtesy of Kim E. Boonekamp and Michael Boutros. Scale bars (a,c) $20\,\mu\mathrm{m}$.
Recently, intestinal organoids have emerged as a model system for curved epithelial sheets. Indeed partially extruded cells can be readily observed in such systems, cf. Fig. 10(a). Moreover one often observes varying cellular opening angles, compare Fig. 10(c), which have been neglected and not found in standard formulations of the VM [36, 37, 38]. As shown in Fig. 10(b) and (d), the BVM can result in very similar shapes to the ones observed for the organoids, if the apico-basal tensions are reduced at pentagonal defects, as predicted by our theory.
@@
Our model makes experimental predictions concerning a facilitated extrusion at pentagonal cells and a curvature dependence of the necessary tension decrease for extrusion. These predictions could be tested in the future in systems such as intestinal organoids in scaffolds [39], where the intestinal epithelium is reconstituted by considering an intestinal monolayer on a 3D substrate with similary shape as found in the intestine. In particular, in this approach cell extrusion has indeed been found to occur at the villus tips [39] and it could be used in the future to investigate our proposed dependence on the radius of background curvature. Intestinal organoids usually display an inverted sign of curvature compared to the villus domain in the intestine with the apical side pointing toward the lumen. Polarity reversed organoids, however, have the opposite apico-basal orientation and thus the correct curvature. They thus constitute another possible test system to study connections of single-cell shape and tissue curvature [40, 41], see an example image in Fig. 10(c).
@@
Our analytical model does not fully describe the morphology of partially extruded cells, because we neglected the relaxation of tilt in the neighboring cells. To better capture such intermediate morphologies of cell extrusion, as observed numerically in the BVM, and thus also obtain better predictions of stability, it is necessary to find a continuum description for these states. One possibility would be to introduce a continuous tilt field, to allow for deviations in the lateral configurations. It has been introduced before in the context of lipid membranes [42, 43], constituting a higher order correction in the continuum theory of bending. Together with the framework of a discrete cell in a mean-field tissue, which we presented here, it can serve as a starting point for further investigations of the coupling of single-cell and tissue-scale mechanics.
@@ Acknowledgements.
We thank Kim E. Boonekamp and Michael Boutros for helpful discussions and for providing the microscopy data from Fig. 10. This research was conducted within the Max Planck School Matter to Life supported by the German Federal Ministry of Education and Research (BMBF) in collaboration with the Max Planck Society. USS is member of the Interdisciplinary Center for Scientific Computing (IWR) at Heidelberg.
@@ Appendix A Description of OrganoidChaste
where we assumed a mobility of one. This equation of motion is solved using Euler stepping. Volume conservation is achieved by projecting the forces on the manifold of volume conserving forces with an additional correction step [32].
@@
Cells can perform neighbor exchanges, so-called T1 transitions, where we first form a protorosette and then in a next step create the new interface, following the approach suggested in Ref. [26]. We perform these randomly with a probability given by a corresponding transition rate on the entire epithelium. This has been found to fluidize the 3D VM, allowing for a viscous relaxation of stresses [26]. In the minimization of the spherical epithelia a decreasing transition rate is taken, similarly to simulated annealing.
@@
In order to solve the bubbly vertex model the configuration which is obtained in OrganoidChaste is exported to Surface Evolver [32], where the same energy, Eq. (1), is considered. However, the interpolation points of the triangulation are not longer averaged, allowing for interfacial curvature. If an interfacial bending rigidity is considered, then the mean curvature in Surface Evolver is calculated at triangulation points of the interfaces, which do not belong to the edges of three intersecting interfaces. This leads to small deviations at the corner triangles due to the neglection of a bending energy across the corresponding interface internal edge (cf. Fig. 9(c)), which decreases as the triangulation resolution is increased.
Appendix B Continuum parameters for the mean field tissue

### v2

@@ IV Discussion
Here we have introduced the bubbly vertex model (BVM) for multicellular epithelial monolayers as a generalization of the standard vertex model (VM). In contrast to the VM, the BVM allows for curved interfaces, which are known to occur in flat epithelia and even more so in curved ones. Because curved epithelial sheets naturally have topological defects due to Euler’s polyhedron theorem, we asked if cell bulging and extrusion might occur naturally at such defects. By comparing the BVM with the VM, we found that allowing for cell interface curvature dramatically lowers the conical buckling threshold around such defects and yields a bulging instability.
To explore the significance of these theoretical predictions, we have performed experiments with spherical mouse colon organoids, which are an appropriate model system for the spherical epithelia described by our theory. Because extrusions are hard to capture due to their speed, we focused on cell shape at topological defects. We found that indeed cell shape is different at pentagonal cells, but differently from the predictions of the BVM, the opening angles were decreased rather than increased. This suggests the existence of some mechanisms for stabilization. Very recently, basal contractility has been reported as such a mechanism [8]. Leveraging a force inference pipeline that we have developed earlier [39], we indeed found an increased basal interfacial tension. Perturbation of the actomyosin cytoskeleton increased the bulging, showing its involvement in actively stabilizing cell shape. We found pentagonal mid-plane areas to be smaller and negatively correlated with basal tensions, potentially facilitating stabilization via mechanosensing. In simulations we found this strategy of cell-specific tension variability to yield full control over cell shapes. We also showed that tissue-scale forces must be at play, in particular pressure, which we estimate to be of the order of $440\,\mathrm{Pa}$. This is one order of magnitude smaller than reported for epithelia compressing hydrogel spheres [50], but likely larger than in mouse small intestine organoids, which use lower pressures as a mechanism for budding [48].
@@
In order to better understand the physical mechanism behind the instability, we developed a theory of a discrete defect cell in a mean-field tissue. This approach provides a mechanistic model for extrusion-like cellular shapes and it suggests that lowering the apico-basal tension in the cell to be extruded, and thus lowering the contractile forces and the mechanical stability of its cortex, provides a mechanism to extrude cells. This is consistent with recent experiments, showing extrusion upon cortical ablation in intestinal organoids and tension-related extrusion in intestinal villi [8]. Line tensions, as they could arise from supracellular actin structures, facilitate extrusion, but do not allow for the extrusion of partially extruded or fully bulged cells as they are observed in some experimental cases.
@@
Our model makes experimental predictions concerning a facilitated extrusion at pentagonal cells and a curvature dependence of the necessary tension decrease for extrusion. These predictions could be tested in the future in systems such as intestinal organoids in scaffolds [51], where the intestinal epithelium is reconstituted by considering an intestinal monolayer on a 3D substrate with similar shape as found in the intestine. In particular, in this approach cell extrusion has indeed been found to occur at the villus tips [51]. It has recently been used to identify tissue tension and cytoskeletal stability to play a role in extrusion [8] and could in the future be used to investigate our proposed dependence on the radius of background curvature. Recently, also two-dimensional systems to study extrusion have become available [52] and instabilities could be investigated in such a setting as well. Intestinal organoids usually display an inverted sign of curvature compared to the villus domain in the intestine with the apical side pointing toward the lumen. Polarity reversed organoids, however, have the opposite apico-basal orientation and thus the correct curvature. They thus constitute another possible test system to study connections of single-cell shape and tissue curvature [53, 54]. Although hard to implement in practice, live cell microscopy might be used to capture extrusion events in more detail and to correlate them with the topological environment. Then it would be interesting to also extend our theoretical framework to include more dynamic processes, such as cellular rearrangements and tissue fluidity, to achieve a more realistic representation of morphology in dynamic tissues. Further, more systematic investigations of different extrusion mechanisms, such as basal protrusions (lamellipodia) [52], present another interesting avenue for future work.
@@
Our analytical model does not fully describe the morphology of partially extruded cells, because we neglected the relaxation of tilt in the neighboring cells, i.e., a non-vanishing angle between the tissue-mid-plane normal and lateral interfaces. To better capture such intermediate morphologies of cell extrusion, as observed numerically in the BVM, and thus also obtain better predictions of stability, it is necessary to find a continuum description for these states. One possibility would be to introduce a continuous tilt field, to allow for deviations in the lateral configurations. It has been introduced before in the context of lipid membranes [55, 56]. Tilt has been linked to scutoid cell shapes, which should be investigated in this context as well [57, 58]. Together with the framework of a discrete cell in a mean-field tissue, which we presented here, it can serve as a starting point for further investigations of the coupling of single-cell and tissue-scale mechanics.
@@ Acknowledgements.
OMD, MB and USS acknowledge support by the Max Planck School Matter to Life, with funding by the German Federal Ministry of Education and Research (BMBF), the Dieter Schwarz Foundation and the Max Planck Society. MB and USS acknowledge support by the cluster of excellence 3DMM2O (EXC 2082/1-390761711 and EXC 2082/2-390761711) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). The authors acknowledge the data storage service SDS@hd supported by the Ministry of Science, Research and the Arts Baden-Württemberg (MWK) and the DFG through grant INST 35/1503-1 FUGG. OMD thanks Edouard Hannezo for valuable discussions. USS is member of the Interdisciplinary Center for Scientific Computing (IWR) at Heidelberg.
@@ Appendix A Description of OrganoidChaste
where we assumed a mobility of one. This equation of motion is solved using Euler stepping. Volume conservation is achieved by projecting the forces on the space of volume conserving forces via a Lagrange multiplier, corresponding to pressures, with an additional correction step. This implementation is identical in our VM code and in the subsequent BVM minimization in Surface Evolver [36] described below.
@@
Cells can perform neighbor exchanges, so-called T1 transitions, where we first form a protorosette (4-way one-dimensional junction) and then in a next step create the new interface, following the approach suggested in Ref. [30]. We perform these randomly with a probability given by a corresponding transition rate on the entire epithelium. This has been found to fluidize the 3D VM, allowing for a viscous relaxation of stresses [30]. In the minimization of the spherical epithelia a decreasing transition rate is taken, similarly to simulated annealing. The resulting organoid is taken as starting point for a comparative minimization in the VM and the BVM with the same underlying topology.
@@
In order to solve the bubbly vertex model the configuration which is obtained in OrganoidChaste is exported to Surface Evolver [36], where the same energy, Eq. (1), is considered. However, the interpolation points of the triangulation are not longer averaged, allowing for interfacial curvature. Despite this, the model and the minimization through overdamped relaxation are identical to the VM if no additional effects are added; in particular the energy is unchanged, Eq. (1). If an interfacial bending rigidity is considered, then the mean curvature in Surface Evolver is calculated at triangulation points of the interfaces, which do not belong to the edges of three intersecting interfaces. This leads to small deviations at the corner triangles due to the neglection of a bending energy across the corresponding interface internal edge (cf. Fig. 4(c)), which decreases as the triangulation resolution is increased.
Simulations of the VM were run on the timescale of minutes on a regular computer. Minimizations in Surface Evolver for large systems were run on a high performance computing cluster on AMD Epyc 7H12 processors, with a minimization running on one $2.6\,\mathrm{GHz}$ core for a time from a few hours to up to three days for very large systems, cf. Fig. 5(c).
Appendix B Experimental methods
Mouse colon (mCol) organoids were generated from a male C57BL/6J mouse as described in Ref. [38]. To maintain organoid cultures, mCol organoids were enzymatically and mechanically dissociated into single cells using TrypLE (Gibco). Cells were resuspended in Basement Membrane Extract Type R1 (BME-R1) (R&D Systems), embedded in $10\,\mu\text{L}$ domes in wells of 12-well plates (Corning), resulting in 5 domes per well. Organoids were cultivated in WENR medium prepared by the addition of $1\times\mathrm{B-27}$ supplement (Gibco), $1.25\,\text{mM}$ n-Acetyl Cysteine (Sigma-Aldrich), $100\,\mathrm{mg}\,\mathrm{mL}^{-1}$ Primocin (InvivoGen), $50\,\mathrm{ng}\,\mathrm{mL}^{-1}$ EGF (Peprotech), $10\%$ (v/v) Noggin Conditioned Medium, $20\%$ (v/v) R-spondin Conditioned Medium, $10\,\mu\mathrm{M}$ Y-27632 (TargetMol) and $0.5\,\mathrm{nM}$ Wnt Surrogate (Gibco) into AD3+ basal medium, which consists of Advanced DMEM/F-12 (Gibco) supplemented with $10\,\mathrm{mM}$ HEPES (Sigma‑Aldrich), $2\,\mathrm{mM}$ GlutaMAX (Gibco), and $10\,\mathrm{mM}$ Penicillin/Streptomycin (Sigma‑Aldrich). Organoids were cultured at $37^{\circ}\mathrm{C}$ in a humidified atmosphere with $5\%$ $\mathrm{CO}_{2}$ (Binder). The culture medium was regularly changed every two to three days, and passaging was performed at a 1:6 ratio approximately every five to seven days.
mCol organoids were dissociated into single cells as described above, and seeded at a density of $1000\,\mathrm{cells}\,\mu\mathrm{L}^{-1}$ BME-R1. The organoids were embedded in 10 $\mu\mathrm{L}$ BME-1 domes in 1 well of 8-well $\mu$-slide (ibidi), and cultured at $37^{\circ}\mathrm{C}$ in a humidified atmosphere with $5\%$ $\mathrm{CO}_{2}$. The culture medium was changed on the second day of culture. For Latrunculin-A perturbation, organoids were treated for 5 hours with $0.5\,\mu\mathrm{M}$ Latrunculin-A (Cayman Chemical). In the corresponding paired experiment the wildtype organoids were left in WENR medium. Following treatment, organoids were fixed ($4\%$ formaldehyde) for 30 minutes at room temperature. $0.5\mu\mathrm{M}$ Latrunculin-A was added to the fixation solution for organoids treated with Latrunculin-A. For immunofluorescence staining, organoids were permeabilized in $0.5\%$ Triton-X100 in PBS (Sigma-Aldrich) for 10 minutes at room temperature. Organoids were subsequently incubated in blocking buffer ($4\%$ BSA, $0.1\%$ Triton X-100, and $0.1\%$ Tween-20 in PBS) for 1 hour at room temperature. The organoids were then incubated overnight at $4^{\circ}\mathrm{C}$ with primary antibody (Mouse anti-Beta-Catenin, BD Biosciences, 1:500). Organoids were washed with $0.1\%$ Tween-20 in PBS for 3 times before secondary antibody incubation for 1 hour at room temperature (Goat-anti-Mouse-AF488, Life Technologies, 1:500). Co-staining for F-actin (Phalloidin-TRITC, Sigma, $50\,\mathrm{ng}\,\mathrm{mL}^{-1}$) and DNA (Hoechst, Life Technologies, $1\,\mu\mathrm{g}\,\mathrm{mL}^{-1}$) was performed simultaneously. After three additional washes with $0.1\%$ Tween-20 in PBS, the organoids were immersed in clearing solution ($\mathrm{2.5}\,\mathrm{M}$ fructose in $60\%$ glycerol).
Images were acquired on a Nikon AX laser scanning confocal microscope using a $25\times$ silicone immersion objective, with an XY pixel size of $0.25\,\mu\mathrm{m}$ and a z-step size of $0.5\,\mu\mathrm{m}$. Images were deconvolved using the automatic 3D deconvolution feature in NIS Elements software Version 5.42.03.
Appendix C Image segmentation
For image segmentation the images were preprocessed by interleaving the image stack with averaged images to obtain cubic voxels. Further, the pixel intensity was linearly corrected as a function of imaging depth for some organoids, to correct for intensity loss. Finally, in some organoids gamma correction was performed to increase the image contrast. Using Cellpose 2.0 [59, 60] and ilastik [61], using the models trained specifically for organoid imaging data with mouse small intestine organoid data in Ref. [39], we segmented the spherical mouse colon organoids.
Using the force inference pipeline we have developed [39], we then performed three-dimensional reconstruction of the interfaces to obtain normal vectors at the tri-cellular junctions and subsequent force inference. For this we assume constant surface tensions across cell-cell, cell-lumen or cell-outside interfaces. This implies the force balance equation at a junction (cf. Fig. 2(c))
$aligned\Gamma_{1}+\Gamma_{2}\cos(\theta_{12})+\Gamma_{3}\cos(\theta_{13})&=0\\ \Gamma_{2}\sin(\theta_{12})-\Gamma_{3}\sin(\theta_{13})&=0,$ (13)
where the three interfaces haves tensions $\Gamma_{1,2,3}$, respectively, and the angle between interface $i$ and $j$ is denoted with $\theta_{ij}$. Including all interfaces and all junctions yields a large overdetermined system of equations, where the angles can be used to infer the underlying (relative) surface tensions $\Gamma$ [39], with the average tension over all interfaces set to one. As in Ref. [39] interfaces which could be associated with segmentation errors were excluded from the analysis.
Both in the BVM and in the experimental data we determine the opening angle of a lateral interface by calculating the angle between the apico-basal axis (the vector pointing from the apical (luminal) interface center of mass to the basal (outer) one) and the inward pointing lateral interface normal $\varphi$ via $\delta/2=\pi/2-\varphi$. For each cell these lateral opening angles are averaged. For this calculation interfaces are represented as point clouds. In the BVM we consider the triangulation points. For the experimental data we obtain the points by considering voxels that have two cell indices in their 19-point stencil (next-to-nearest neighbor voxels) and taking their positions in voxel space. We determine the lateral normals as the direction with smallest variance via a principal component analysis [39].
To determine the number of neighbors for each cell, we consider the basal (outer) side, as we found segmentation errors to usually occur on the inner apical side. In the segmentation we sometimes found inclusions of small cells due to segmentation issues [39]. We thus neglected cells with a volume below $1000$ voxels ($15.625\,\mu\mathrm{m}^{3}$) and then determined tri-interfacial junctions, again via the 19-point stencil. For each basal interface the number of neighboring junctions, i.e., polygonal edges, was determined as the number of neighbors. Cells in which segmentation errors were identified (as in the force inference step) are then neglected in the statistics. Their basal sides, however, have contributed to the neighbor determination thus not leading to an error from this correction.
Appendix D Estimating luminal pressure
To estimate luminal pressures we consider the vertex model with flat interfaces and its coarse-grained nonlinear stretching energy density [19]
$e_{\mathrm{stretch}}=\frac{\Gamma_{\mathrm{a}}+\Gamma_{\mathrm{b}}}{\Gamma_{\mathrm{l}}}\,\Gamma_{0}\left(\lambda^{2}+2\frac{1}{\lambda}-3\right),$ (14)
with average tension $\Gamma_{0}=5\,\mathrm{nN}\,\mu\mathrm{m}^{-1}$ and principal stretches in the spherical shell $\lambda=\lambda_{\phi}=\lambda_{\theta}$, which describe an elongation $\ell^{\prime}=(1+\lambda)\ell$ for an in-plane length $\ell$ which is stretched to $\ell^{\prime}$. One can show that bending energy contributions constitute higher order corrections which can be neglected.
To obtain the pressure, we consider
$P\,\mathrm{d}[V(R)]=\mathrm{d}[e_{\mathrm{stretch}}(R)A_{0}],$ (15)
with pressure P, spherical volume $V$ and undeformed spherical area $A_{0}$; the latter two both depend on the pressured sphere radius $R$. For a stretched sphere the principal stretch is $\lambda=R/R_{0}$ with undeformed sphere radius $R_{0}$. We can insert the volume and area of a sphere to obtain
$P\,4\pi R^{2}\,\mathrm{d}R=4\pi\,\frac{\Gamma_{\mathrm{a}}+\Gamma_{\mathrm{b}}}{\Gamma_{\mathrm{l}}}\,\Gamma_{0}\left(2R-2\frac{R_{0}^{3}}{R^{2}}\right)\mathrm{d}R,$ (16)
which implies for the pressure
$P=\frac{\Gamma_{\mathrm{a}}+\Gamma_{\mathrm{b}}}{\Gamma_{\mathrm{l}}}\,\Gamma_{0}\left(\frac{2}{R}-2\frac{R_{0}^{3}}{R^{4}}\right).$ (17)
Note that the first term is the Young-Laplace law which is modified by the second term stemming from the initial shell radius. For the shell in Fig. 2(a) with $\Gamma_{\mathrm{a}}\approx 1.02$, $\Gamma_{\mathrm{b}}\approx 1.06$, $\Gamma_{\mathrm{l}}\approx 0.97$, and $R\approx 22\,\mu\mathrm{m}$ we have $R_{0}=\left(\psi_{\mathrm{exp}}/\psi_{\mathrm{theo}}\right)^{1/3}R\approx 18\,\mu\mathrm{m}$, which implies $P\approx 440\,\mathrm{Pa}$.
Appendix E Effect of line tension
Figure 10: Effect of basal (outer) line tensions in all cells on the opening angles of defect cells. (a) Opening angles as function of the dimensionless line tension $\tau$. Individual red lines mark single pentagonal cells to show single-cell dependency of inward bulging. Symbols show means with shaded regions indicating $5\%$ and $95\%$ quantiles for hexagons in black and heptagons in blue. (b) Spherical shell as in Fig. 1 for $\tau=0.05$.
The existence of line tensions has been demonstrated in small intestine organoids [39] and could in principle stabilize cell shapes basally. To consider such a process in our model, we add a line tension at all outer basal interface boundaries, which effectively constitutes a contractile contribution along the basal interface edges. Fig. 10 displays the opening angles of BVM simulations with basal (outer) line tensions for all cells, where we see cell-specific variability in switching opening angles through an unbulged intermediate step. The cell-specific switching again suggests that cell-to-cell variability in tensions is necessary.
Besides a reduction of apico-basal interfacial tension, another possible mechanism for extrusion are contractile actin structures pushing cells outward [3]. To introduce this, we consider a line tension on the (luminal) apical side. For a tension $\tau$ we thus have an energy at the defect
$E_{\mathrm{line}}=n\,\tau\,a,$ (18)
which is an additional $a$-dependent term contributing to the total energy. It leads to a tilting of the energy landscape in configuration space, because in the case of curvature screening no $R_{0}$ minimization is performed. Fig. 11 displays such energy landscapes for pentagons ($n=5$) for different line tensions $\tau$. We do see that the tilting can destabilize bulged states. However, since at $a=0$ the line tension does not contribute anymore, an energy barrier in the case of a transitioning from a bulged state with $a=0$ to the extruded case is not changed by the line tension. In light of partially extruded cells, such as in Fig. 8(g), this implies that an additional tuning of interfacial tensions (either lateral or apico-basal) is necessary to achieve full extrusion.
Appendix F Details on mean-field tissue parameterization
The deformation energy for a tissue parametrized as in Fig. 6 comprises bending and stretching components, derived in the following. For the cap segment we consider the radial deflection function $f$ with rotational symmetry
$f(r)=R_{0}-\sqrt{R_{0}^{2}-(\rho+r)^{2}},$ (19)
$e_{\mathrm{bending}}=\frac{\kappa}{2}H^{2}+\kappa_{\mathrm{G}}K,$ (20)
$aligned H(r)&=\frac{1}{r}\partial_{r}r\left(\frac{\partial_{r}f}{\sqrt{1+(\partial_{r}f)^{2}}}\right)=\frac{2}{R_{0}}+\frac{\rho}{rR_{0}},\\ K(r)&=\frac{\frac{1}{r}(\partial_{r}f)(\partial_{r}^{2}f)}{\left(1+(\partial_{r}f)^{2}\right)^{2}}=\frac{\rho+r}{rR_{0}^{2}}.$ (21)
{multline}Ebendingmeanκπ
= - 4R0 R_0^2 - (ρ+r)^2 - ρ2R02-ρ2 ×
artanh ( R02-ρ( ρ+ r )(R02- ρ2)1/2( R02- (ρ+r)2)1/2) —_r=r_0^r=R_B. Note that for $\rho=r_{0}=0$ and small buckling radius $R_{\mathrm{B}}$, i.e. small $|\nabla f|$ in the outer cone segment, the Taylor approximation in $R_{\mathrm{B}}$ scales quadratically, consistent with the approximation of the total energy with a flat disc as proposed by Seung and Nelson [23] for the flattened tip for the deformed shape around the defect. The first term in Eq. (F) arises from the spherical contribution of the mean curvature bending energy, cf. corresponding first term in the mean curvature in Eq. (21). The second term, which is proportional to the squared radial offset $\rho^{2}$, serves as a correction for reducing the radius of the circular segment in cylindrical coordinates, as we decrease the radius of curvature in the azimuthal direction via this offset. This term stems from the quadratic contribution of $\rho^{2}$ from the squared mean curvature, Eq. (21), in the bending energy.
with tissue opening angle $\gamma/2$, cf. Fig. 6(a).
$E_{\mathrm{cone}}=\frac{f_{1}}{(1+f_{1}^{2})^{1/2}}\pi\kappa\log(R/R_{\mathrm{B}}),$ (23)
with the total radius $R$, the buckling radius $R_{\mathrm{B}}$, and the cone slope
$f_{1}^{\mathrm{cone}}=\left[\frac{1}{(1-s/2\pi)^{2}}-1\right]^{1/2}.$ (24)
The stretching energy density reads
$aligned\varepsilon_{rr}&=w^{\prime}(r)-1+\frac{(f^{\prime}(r))^{2}}{2},\\ \varepsilon_{\theta\theta}(r)&=\frac{w(r)-r}{r}+q_{s},\\ \varepsilon_{r\theta}&=0,$ (26)
We now want to take a different approach, however, and consider a ring of material with inner and outer radii $r_{0}$ and $R_{\mathrm{B}}$, deflected with a known fixed slope $f^{\prime}=f_{1}$. Introducing $\Delta s=q_{s}-{f_{1}}^{2}/2$, Eq. (27) is solved by
$w(r)=\frac{\mu}{2\mu+\lambda}\Delta sr\log(r)+\frac{k_{1}}{r}+k_{2}r,$ (28)
$\Delta s=q_{s}-\left(1-\frac{1}{\sqrt{1+{f_{1}}^{2}}}\right).$ (29)
As boundary conditions we assume vanishing radial stresses at the inner and outer radii, i.e. at $r=r_{0}$ and $r=R_{\mathrm{B}}$
This boundary condition assumes that no radial (and thus total compressive) stresses act on the boundary of the cap segment and the conical part of the surrounding tissue. Comparing the cell height in VM simulations to the theoretical mean-field tissue height serves as an estimate for the compressive in-plane stress. The height difference from the mean-field height (cf. Appendix H) is shown in Fig. 12(a).
Figure 12: Comparison of continuum model with vertex model simulations of icosahedral shells. (a) VM shell with Caspar-Klug size $(20,0)$ and $\Gamma=1.1$ with top view and side view cross-section. Color-coded is cell height difference from the mean-field VM height (cf. Eq. (34)). (b) Total energy of the proposed full defect model as function of defect opening angle $\delta/2$ for different apical/basal surface tensions $\Gamma$. The dotted grey line indicates the angle boundary from the volume constraint. (c) Opening angles $\delta/2$ (top) and apical to basal face area ratios (bottom) for the different models. The defect model (solid) considers a defect in a mean-field tissue with the angle $\delta/2$ corresponding to the minimal energy configuration, cf. (a). The spherical cap model (dashed) corresponds to this energy under the assumption that $\delta/2$ matches the background curvature. The symbols indicate vertex model simulation results for different Caspar-Klug indices $(T_{1},T_{2})$.
$aligned\frac{E_{\mathrm{stretch}}}{\pi Y\left(1+f_{1}^{2}\right)^{1/2}}&=\frac{\Delta s^{2}}{8}\left(R_{\mathrm{B}}^{2}-r_{0}^{2}\right)\\ &\ -\frac{\Delta s^{2}}{2}\ \frac{R_{\mathrm{B}}^{2}r_{0}^{2}}{R_{\mathrm{B}}^{2}-r_{0}^{2}}\log^{2}\left(\frac{R_{\mathrm{B}}}{r_{0}}\right).$ (31)
Finally, we can write down the full energy of the surrounding tissue as a sum of the contributions in Eqs. (F, 22, 23, 31), i.e.
Appendix G Verification of defect cell treatment in icosahedral shells
The corresponding continuum model energy for this vertex model case consists of the core energy for a defect cell without bulging apical/basal membranes, i.e., with $h_{S}^{a}=0=h_{S}^{b}$, Eq. (III.1); the bending energy of both the cap segment, Eqs. (F, 22), and the outer cone region, Eq. (23); and the stretching energy in the cap, Eq. (31). For the resulting total energy $E_{\mathrm{tot}}$ we assume a large outer radius of the total disc $R$, such that area changes from deformations in the cone are negligible, and thus drop the term proportional to $\log(R)$ in Eq. (23), as we look at energy differences with respect to a reference state.
In the derivation of the mean field tissue (cf. Appendix H) lateral interface energies in the energy density are included by counting them half for the respective cell area (to avoid double counting). Here we now introduce a boundary energy term, which subtracts this lateral energy from the mean field tissue and adds half the pentagonal lattice energy of the defect. This lateral energy is present in full and should be considered completely, due to the explicit description of the defect cell. For the lateral contact between the defect and the tissue we thus subtract the lateral energy contribution which is implicitly given in the mean-field tissue boundary (with a circular cut-out) and replace it by the actual lateral energy contribution of the cell (with a pentagonal shape). This energy reads
$E_{\mathrm{bndry}}=\frac{5\tilde{h}}{4}\left(2a+\frac{\tilde{h}}{\eta_{\mathrm{geom}}}\sin(\delta/2)\right)-\pi r_{0}\tilde{h}$ (33)
and it is added to the total energy, i.e. {multline} E_tot = E_core + E^mean_bending + E^Gauss_bending + E_cone
+ E_stretch + E_bndry.
For the elastic parameters and the bending rigidities of the mean-field tissues we use the mean-field values derived for the VM in Ref. [19], i.e. Eqs. (35, 37) in Appendix H.
We found that for every $\delta/2$ and $a$ the energy has a unique minimum as a function of the buckling radius $R_{\mathrm{B}}$, or, equivalently, as a function of the radial cap curvature radius $R_{0}$. We thus minimize the energy with respect to $R_{\mathrm{B}}$ to obtain the total energy as a function of the opening angle $\delta/2$ in the vertex model. Fig. 12(b) depicts the energies for different surface tensions $\Gamma$, where we observe a clear minimum. Note that for large $\Gamma$ the volume constraint yields an upper bound for the possible opening angles. Our approach, the defect model, allows the defect to have an opening angle independently from the background tissue curvature. For comparison we also consider the case of the opening angle following the background curvature, i.e., $\delta/2$ is chosen such that it matches $R_{0}$ with $\rho=0$, which we henceforth denote as the spherical cap model. The offset of the energies is a consequence of the neglection of the $\log(R)$-term and results from the different cell number densities for different $\Gamma$.
Appendix H Continuum parameters for the mean field tissue